# Benefits of Mutual Coupling in Dynamic Metasurface Antennas for Optimizing Wireless Communications – Theory and Experimental Validation

### v1

@@ V-A2 Experimental Setup
We measure the field radiated by the DMA with a broadband double-ridged horn antenna (DRH67) that is mounted on a two-axis linear translation stage. The horn antenna faces the DMA, as seen in Fig. 5. Our choice for the horn antenna is motivated by its low-dispersion broadband matching because our time-gating technique requires an acquisition of the DMA’s radiated field across a broad spectrum, ideally without dispersive distortions from the receiving antenna. Of course, the field captured by the horn antenna must be understood as a weighted integral of the field over parts of the scene rather than the field at a single point, but this does not pose an issue for our present study. Both the DMA and the horn antenna are connected via coaxial cables to a vector network analyzer (Rhode & Schwarz ZVA 67) that measures the transmission coefficient between the two antenna ports for 16,001 linearly spaced frequency points between 15 GHz and 26.5 GHz. We denote this transmission coefficient by $h(f,\mathbf{b})$ because it is the wireless channel between the two antennas’ ports, parametrized by the DMA configuration $\mathbf{b}$.
@@ V-A3 Measurement Procedure
1 for each considered horn antenna position do
2       for $5120<r<$ 512 do
@@ V-C Experimental Results
We first inspect the average absolute magnitudes of the wireless channel, the entries of the Jacobian and the off-diagonal entries of the Hessian in Fig. 6 as a function of $\tau$. It is important to interpret the $\tau$ dependence in light of the average channel impulse response envelope shown on top in Fig. 6. The LOS delay of roughly 3 ns makes sense in sight of the separation of the two antennas and the propagation inside the antennas, and it implies that no signal can be received for $\tau<3\ \mathrm{ns}$ (except for the build-up of the finite-bandwidth LOS tap). Results for $\tau<3\ \mathrm{ns}$ can hence not be interpreted meaningfully, and are measured with a poor signal-to-noise ratio (the wireless channel magnitude is an order of magnitude lower at $\tau=2.75\ \mathrm{ns}$ than at $\tau=3\ \mathrm{ns}$). It is apparent in Fig. 6 that the Jacobian and Hessian entries are, respectively, one and two order of magnitude weaker than the wireless channel. In the limit $\tau\rightarrow\infty$, all three quantities converge to their values without time gating that are indicated by horizontal dashed lines. The fact that the wireless channel’s absolute magnitude varies with $\tau$ emphasizes the importance of studying the normalized Jacobian and Hessian to make a fair comparison with fixed radiation efficiency, as we do in Fig. 7. Upon careful visual inspection, it is apparent in Fig. 6 that the wireless channel converges quicker to the dashed reference than the Jacobian, which in turn converges much quicker than the Hessian. This observation reveals that the relative DMA radiation pattern sensitivity increases for larger $\tau$ (and, hence, for larger effective DMA mutual coupling strengths), and that the relative non-linearity of the DMA radiation pattern parametrization increases even stronger for larger $\tau$. These trends confirm that the mutual coupling strength governs the trade-off between strength and simplicity of the dependence of the radiation pattern on the DMA configuration.
@@
To reveal these trends more directly, we plot the average relative magnitudes of the Jacobian, $\langle\sigma\rangle$, and Hessian, $\langle\theta\rangle$, in Fig. 7. The average relative sensitivity without time gating (i.e., for $\tau\rightarrow\infty$) is 45.4 % stronger than for time-gating at the LOS delay $\tau=3\ \mathrm{ns}$. In other words, stronger mutual coupling in the DMA results in a substantially larger sensitivity of the radiated field to the DMA configuration. This larger sensitivity translates into massive benefits for wireless engineers to tailor the radiation pattern. Indeed, for the prototypical problem of enhancing the UE channel gain, we observe that the 45.4 % improvement in relative sensitivity translate into 219.0 % improvement in the channel gain enhancement, which rises from 11.6 at $\tau=3\ \mathrm{ns}$ to 25.4 for $\tau\rightarrow\infty$. The flipside of this benefit of stronger mutual coupling is the increased non-linearity of the radiation pattern parametrization. Indeed, the linearity metric $\langle\zeta\rangle$ drops from 32.1 dB at $\tau=3\ \mathrm{ns}$ to 20.7 dB for $\tau\rightarrow\infty$. Correspondingly the average relative magnitude of the off-diagonal entries of the Hessian double from their minimum at $\tau=3.25$ ns to the limit of $\tau\rightarrow\infty$. As explained earlier, the values for $\tau<3$ ns cannot be meaningfully interpreted because they correspond to unphysical delay times. In the case of the Hessian, even the value at $\tau=3$ ns appears abnormal, and only from $\tau=3.25$ ns onwards the trend is clear. This may be due to the fact that the Hessian originates from trajectories within the DMA that have encountered at least two meta-atoms and correspond hence to delay times that are necessarily larger than the LOS delay time of 3 ns. In other words, for the LOS delay time of 3 ns we cannot expect any meaningful measurement of the Hessian.
Figure 7: Experimental results for the trade-off between strength and simplicity of the dependence of $h(\mathbf{b})$ on $\mathbf{b}$, governed by the DMA’s mutual coupling strength. We use the time-gating truncation time $\tau$ to control the DMA’s effective mutual coupling strength in post-processing. The larger $\tau$ is, the stronger is the effective mutual coupling strength (see main text for detailed explanations). The top row reproduces for reference the average impulse response envelope, and the pink vertical line indicates again the LOS delay time. The subsequent rows display the average sensitivity $\langle\sigma\rangle$, the average channel gain enhancement $\langle\eta\rangle$, the average normalized off-diagonal Hessian entries $\langle\theta\rangle$, and the average linearity metric $\langle\zeta\rangle$, for different time-gating truncation times $\tau$. In the limit $\tau\rightarrow\infty$, the curves converge to the horizontal dashed lines obtained without time-gating.
@@ V-D Summary
In this section, we introduced a time-gating technique to control the effective mutual coupling strength in post-processing. We then leveraged this technique to report experimental evidence based on a DMA prototype in the lower part of the K-band. We experimentally measured the Jacobian and Hessian of the radiation pattern. We experimentally validated the trends seen in previous sections, namely that a stronger mutual coupling strength results in a larger sensitivity (which translates into substantial improvements in channel gain enhancement) and a larger Hessian (which translates into a lower linearity metric).
VI Conclusion
@@
To summarize, we have demonstrated that the mutual coupling strength in a DMA governs a fundamental trade-off between strength and simplicity of the dependence of the DMA’s radiation pattern on the DMA’s configuration, based on systematic theoretical, numerical and experimental evidence. On the one hand, the stronger the mutual coupling between the meta-atoms is, the more sensitive the radiation pattern is to the DMA configuration. We demonstrated that this enhanced sensitivity translates into substantial improvements in our ability to optimize the DMA configuration for a desired wireless functionality, considering channel gain enhancement as prototypical problem. On the other hand, stronger mutual coupling makes the mapping from configuration to radiation pattern more non-linear, complicating the accurate modeling and physics-compliant optimization of DMAs for wireless communications.
@@
Our work unveils untapped potential for substantially improved performance in DMA-assisted wireless communications that was overlooked so far because most existing studies on DMAs neglected or deliberately mitigated mutual coupling. The price for reaping the benefits of strong mutual coupling between meta-atoms is the requirement for physics-compliant modeling and optimization. We presented a compact and implementation-agnostic physics-compliant DMA model and are confident that its parameters can be estimated to describe any given DMA prototype, using calibration techniques similar to those already experimentally validated in the RIS context [45, 48, 49, 38].
@@
Looking forward, our findings may encourage wireless practitioners to fundamentally reconsider the current assessment of mutual coupling in DMAs as a vexing nuance. Concretely, we expect DMA architectures that favor strong mutual coupling such as the chaotic-cavity-backed ones considered in our studies to gain traction, and that efficient DMA optimization algorithms capable of coping with strong mutual coupling will be developed. We also mentioned an idea for a new generation of “beyond-diagonal” DMA hardware with tunable mutual coupling which would enable attaining the upper bound on the radiation pattern sensitivity.
@@
Beyond DMAs, we believe that our work has unveiled fundamental links between the mutual coupling strength and the ability to control waves in reconfigurable wave systems that can be transposed to other types of FAS [3, 4, 5, 6, 7] (notably, reconfigurable pixel-based antennas [8, 9, 10, 11, 12, 13]), as well as to RIS [53], wireless localization [35], and wave-based physical neural networks [61].
@@
$\begin{split}\mathbf{j}_{i}=\left[\frac{\partial\mathbf{W}^{-1}}{\partial c_{i}}\right]_{\mathcal{MF}}&=\left[\mathbf{W}^{-1}\right]_{\mathcal{MM}}\mathbf{E}_{ii}\left[\mathbf{W}^{-1}\right]_{\mathcal{MF}}\\ &=\mathbf{\Omega}_{\mathcal{M}\mathcal{M}_{i}}(\mathbf{c})\mathbf{\Omega}_{\mathcal{M}_{i}\mathcal{F}}(\mathbf{c}).\end{split}$ (30)
@@
Using Weyl’s inequality for singular values [63] [64, Theorem 1] yields $|\sigma_{N_{\mathrm{M}}}-\tilde{\sigma}_{N_{\mathrm{M}}}|\leq\|\mathbf{B}\|_{2}$. Then, using the reverse triangle inequality, we bound $\tilde{\sigma}_{N_{\mathrm{M}}}$ as $\tilde{\sigma}_{N_{\mathrm{M}}}\geq\sigma_{N_{\mathrm{M}}}-\|\mathbf{B}\|_{2}$, provided that $\sigma_{N_{\mathrm{M}}}\geq\|\mathbf{B}\|_{2}$ because otherwise we would have the trivial inequality $\tilde{\sigma}_{N_{\mathrm{M}}}\geq 0$. Finally, defining $\tilde{\mu}\triangleq\|\mathbf{B}\|_{2}.\|\mathbf{A}^{-1}\|_{2}=\frac{\|\mathbf{B}\|_{2}}{\sigma_{N_{\mathrm{M}}}}$ and inverting both sides of the previous inequality yields

### v3

@@ V-A2 Experimental Setup
We measure the field radiated by the DMA with a broadband double-ridged horn antenna (DRH67) that is mounted on a two-axis linear translation stage. The horn antenna faces the DMA, as seen in Fig. 5. Our choice for the horn antenna is motivated by its low-dispersion broadband matching because our time-gating technique requires an acquisition of the DMA’s radiated field across a broad spectrum, ideally without dispersive distortions from the receiving antenna. Of course, the field captured by the horn antenna must be understood as a weighted integral of the field over parts of the scene rather than the field at a single point, but this does not pose an issue for our present study. Both the DMA and the horn antenna are connected via coaxial cables to a vector network analyzer (Rohde & Schwarz ZVA67) that measures the transmission coefficient between the two antenna ports for 16,001 linearly spaced frequency points between 15 GHz and 26.5 GHz. [^8] This wideband measurement is required for our time-gating analysis to suppress multipath contributions beyond a desired delay time at the operation frequency. To be clear, our optimization targets a single frequency of operation, and no claim of broadband DMA operation across 11.5 GHz is implied. We denote this transmission coefficient by $h(f,\mathbf{b})$ because it is the wireless channel between the two antennas’ ports, parametrized by the DMA configuration $\mathbf{b}$.
@@ V-A3 Measurement Procedure
1for each considered horn antenna position do
2  for $0<r<512$ do
@@ V-C Experimental Results
We first inspect the average absolute magnitudes of the wireless channel, the entries of the Jacobian and the off-diagonal entries of the Hessian in Fig. 6 as a function of $\tau$. It is important to interpret the $\tau$ dependence in light of the average channel impulse response envelope shown on top in Fig. 6. The LOS delay of roughly 3 ns makes sense in sight of the separation of the two antennas and the propagation inside the antennas, and it implies that no signal can be received for $\tau<3\ \mathrm{ns}$ (except for the build-up of the finite-bandwidth LOS tap). Results for $\tau<3\ \mathrm{ns}$ can hence not be interpreted meaningfully, and are measured with a poor signal-to-noise ratio (the wireless channel magnitude is an order of magnitude lower at $\tau=2.75\ \mathrm{ns}$ than at $\tau=3\ \mathrm{ns}$). It is apparent in Fig. 6 that the Jacobian and Hessian entries are, respectively, one and two orders of magnitude weaker than the wireless channel. In the limit $\tau\rightarrow\infty$, all three quantities converge to their values without time gating that are indicated by horizontal dashed lines. The fact that the wireless channel’s absolute magnitude varies with $\tau$ emphasizes the importance of studying the normalized Jacobian and Hessian to make a fair comparison with fixed radiation efficiency, as we do in Fig. 7. Upon careful visual inspection, it is apparent in Fig. 6 that the wireless channel converges quicker to the dashed reference than the Jacobian, which in turn converges much quicker than the Hessian. This observation reveals that the relative DMA radiation pattern sensitivity increases for larger $\tau$ (and, hence, for larger effective DMA MC strengths), and that the relative non-linearity of the DMA radiation pattern parametrization increases even stronger for larger $\tau$. These trends confirm that the MC strength governs the trade-off between strength and simplicity of the dependence of the radiation pattern on the DMA configuration.
@@
To reveal these trends more directly, we plot the average relative magnitudes of the Jacobian, $\langle\sigma\rangle$, and Hessian, $\langle\theta\rangle$, in Fig. 7. The average relative sensitivity without time gating (i.e., for $\tau\rightarrow\infty$) is 45.4 % stronger than for time-gating at the LOS delay $\tau=3\ \mathrm{ns}$. In other words, stronger MC in the DMA results in a substantially larger sensitivity of the radiated field to the DMA configuration. This larger sensitivity translates into massive benefits for wireless engineers to tailor the radiation pattern. Indeed, for the prototypical problem of enhancing the UE channel gain, we observe that the 45.4 % improvement in relative sensitivity translate into 219.0 % improvement in the channel gain enhancement, which rises from 11.6 at $\tau=3\ \mathrm{ns}$ to 25.4 for $\tau\rightarrow\infty$. The flipside of this benefit of stronger MC is the increased non-linearity of the radiation pattern parametrization. Indeed, the linearity metric $\langle\zeta\rangle$ drops from 32.1 dB at $\tau=3\ \mathrm{ns}$ to 20.7 dB for $\tau\rightarrow\infty$. Correspondingly the average relative magnitude of the off-diagonal entries of the Hessian double from their minimum at $\tau=3.25$ ns to the limit of $\tau\rightarrow\infty$. As explained earlier, the values for $\tau<3$ ns cannot be meaningfully interpreted because they correspond to unphysical delay times. In the case of the Hessian, even the value at $\tau=3$ ns appears abnormal, and only from $\tau=3.25$ ns onwards the trend is clear. This may be due to the fact that the Hessian originates from trajectories within the DMA that have encountered at least two meta-atoms and correspond hence to delay times that are necessarily larger than the LOS delay time of 3 ns. In other words, for the LOS delay time of 3 ns we cannot expect any meaningful measurement of the Hessian.
Figure 7: Experimental results for the trade-off between strength and simplicity of the dependence of $h(\mathbf{b})$ on $\mathbf{b}$, governed by the DMA’s MC strength. We use the time-gating truncation time $\tau$ to control the DMA’s effective MC strength in post-processing. The larger $\tau$ is, the stronger is the effective MC strength (see main text for detailed explanations). The top row reproduces for reference the average impulse response envelope, and the pink vertical line indicates again the LOS delay time. The subsequent rows display the average sensitivity $\langle\sigma\rangle$, the average channel gain enhancement $\langle\eta\rangle$, the average normalized off-diagonal Hessian entries $\langle\theta\rangle$, and the average linearity metric $\langle\zeta\rangle$, for different time-gating truncation times $\tau$. In the limit $\tau\rightarrow\infty$, the curves converge to the horizontal dashed lines obtained without time-gating.
@@ V-D Summary
In this section, we introduced a time-gating technique to control the effective MC strength in post-processing. We then leveraged this technique to report experimental evidence based on a DMA prototype in the lower part of the K-band. We experimentally measured the Jacobian and Hessian of the radiation pattern. We experimentally validated the trends seen in previous sections, namely that a stronger MC strength results in a larger sensitivity (which translates into substantial improvements in channel gain enhancement) and a larger Hessian (which translates into a lower linearity metric).
VI Discussion
The goal of this paper is to evidence that strong MC enhances the sensitivity of the radiation pattern to the DMA configuration, and that an enhanced sensitivity improves the ability to tailor the radiation pattern to a desired functionality. Consequently, we argue that strong MC should be embraced to maximize the radiation pattern tunability with a given number of tunable elements. In contrast to our argument, the literature contains some works that actively seek to limit (rather than to maximize) the sensitivity [66, 67]. To the best of our understanding, there is no evidence of the connection between MC strength, sensitivity and tunability in these works; instead, their motivation to limit the sensitivity originates from considerations of vulnerability to fabrication inaccuracy and operation bandwidth. In this discussion, we address these two concerns.
VI-A Vulnerability to fabrication inaccuracies
Most schemes for operating reconfigurable antennas in experiments (including [67]) rely on a forward model of the antenna whose parameters are known analytically or obtained in numerical simulations based on the known antenna design. Naturally, such schemes are vulnerable to fabrication inaccuracies. The more sensitive the radiation pattern is to the DMA configuration, the more sensitive it is also to inaccuracies in the model parameters. In such schemes, the vulnerability to fabrication inaccuracies raised in [66, 67] is thus a valid concern.
Our vision for operating reconfigurable antennas in experiments is different. We envision that a compact model is calibrated in situ, i.e., its parameters are estimated for an experimentally given, fabricated device. Thereby, the concern about vulnerability to fabrication inaccuracies inherently vanishes. We are confident that this approach is feasible given our recent works on the conceptually related problem of estimating model parameters for RIS-parametrized wireless channels [47, 50, 51, 38, 68, 69]. For the purpose of experimental parameter estimation, the model should be as compact as possible to avoid over-parametrization resulting in avoidable ambiguities and avoidable computational complexity. Our model formulation in Sec. II is maximally compact because it collapses architecture-specific details of the coupling mechanisms between feeds and meta-atoms into a background coupling operator; in any case, the details of these coupling mechanisms could not be identified unambiguously based on experimental measurements. However, the algorithmic developments to transpose the parameter estimation approaches from [47, 50, 51, 38, 68, 69] in the realm of RIS to DMAs are beyond the scope of this paper and are thus deferred to future work. We emphasize that all required experimental results for this paper are obtained in a model-agnostic manner, so that this paper did not require the development of parameter-estimation algorithms.
VI-B Operation bandwidth
Because of the fundamental connection between dwell time and sensitivity [35], an enhanced sensitivity is also expected to result in a more rapid spectral decorrelation of the radiation pattern. In other words, if the DMA configuration is optimized to achieve a desired functionality at a single frequency, the interval of frequencies around the optimized frequency for which the desired functionality is also roughly achieved is narrower. This effect can indeed be observed in our experimental data, as documented for selected examples in Fig. 8 and summarized on average over all examples in Table I. The larger the truncation time $\tau$ is (which is our proxy for the MC strength), the narrower is the peak around the targeted frequency in the optimized spectrum of the end-to-end channel gain.
Figure 8: Selected examples of end-to-end channel spectra for $f_{0}=19$ GHz for four different choices of $\tau$. (blue: optimized; gray: random; black: average over random)
TABLE I: Experimentally evaluated FWHM bandwidth of the optimized end-to-end channel, averaged over individual optimizations for each receiver position and operating frequency.
$\tau$ [ns] 2.5 2.75 3.00 3.25 3.50 4.00 4.50 $\infty$
BW [GHz] N/A N/A 4.00 2.14 1.46 0.80 0.58 0.35
However, even for the case of $\tau\rightarrow\infty$, we still have an average peak width of 0.35 GHz which can comfortably accommodate at least one wireless link. We emphasize that the relatively narrow bandwidth does not preclude operations of our DMA across a much wider frequency range. On the one hand, we can sequentially switch between DMA configurations optimized for different central operating frequencies. On the other hand, we can optimize the DMA configuration to simultaneously serve multiple carrier frequencies (by trading off channel gain with bandwidth).
Finally, we note that the spectral selectivity associated with strong MC can actually be a desirable feature enabling a frugal use of spectral resources as well as interference rejection. In addition, applications like computational imaging and wireless power transfer often target single-frequency operation to alleviate the cost of the required RF chains.
Altogether, while there is a clear trade-off between MC strength and operation bandwidth, we have shown that the largest conceivable MC strength in our DMA prototype still offers sufficient bandwidth for a typical wireless link. While our optimizations in the present paper were focused on single-frequency optimizations, future work will explore the optimization of the DMA configuration for desired functionalities across multiple frequency points.
VII Conclusion
@@
To summarize, we have demonstrated that a DMA’s MC strength governs a fundamental trade-off between strength and simplicity of the dependence of the DMA’s radiation pattern on the DMA’s configuration, based on systematic theoretical, numerical and experimental evidence. On the one hand, the stronger the MC between the meta-atoms is, the more sensitive the radiation pattern is to the DMA configuration. We demonstrated that this enhanced sensitivity translates into substantial improvements in our ability to optimize the DMA configuration for a desired wireless functionality, considering channel gain enhancement as prototypical problem. On the other hand, stronger MC makes the mapping from configuration to radiation pattern more non-linear, complicating the accurate modeling and physics-compliant optimization of DMAs for wireless communications.
@@
Our work unveils untapped potential for substantially improved performance in DMA-assisted wireless communications that was overlooked so far because most existing studies on DMAs neglected or deliberately mitigated MC. The price for reaping the benefits of strong MC between meta-atoms is the requirement for physics-compliant modeling and optimization. We presented a compact and implementation-agnostic physics-compliant DMA model and are confident that its parameters can be estimated to describe any given DMA prototype, using calibration techniques similar to those already experimentally validated in the RIS context [47, 50, 51, 38, 68, 69].
@@
Looking forward, our findings may encourage wireless practitioners to fundamentally reconsider the current assessment of MC in DMAs as a vexing nuance. Concretely, we expect DMA architectures that favor strong MC such as the chaotic-cavity-backed ones considered in our studies to gain traction, and that efficient DMA optimization algorithms capable of coping with strong MC will be developed. We also mentioned an idea for a new generation of “beyond-diagonal” DMA (BD-DMA) hardware with tunable MC which can enable attaining the upper bound on the radiation pattern sensitivity; an initial exploration of performance gains with BD-DMAs can be found in our recent follow-up work [57].
@@
We expect that the unveiled fundamental dependence of the ability to control the transfer function of a reconfigurable wave system on the MC strength generalizes to ESPAR antennas [3, 4, 5, 6, 7, 8, 9], reconfigurable pixel-based antennas [10, 11, 12], RIS [56, 58, 59], wireless localization [35], and wave-domain physical neural networks [70].
@@
$\begin{split}\mathbf{j}^{\mathrm{p}}_{i}=\left[\frac{\partial\mathbf{W}^{-1}}{\partial c_{i}}\right]_{\mathcal{MF}}e^{\mathrm{inc}}&=\left[\mathbf{W}^{-1}\right]_{\mathcal{MM}}\mathbf{E}_{ii}\left[\mathbf{W}^{-1}\right]_{\mathcal{MF}}e^{\mathrm{inc}}\\ &=\mathbf{\Omega}_{\mathcal{M}\mathcal{M}_{i}}(\mathbf{c})\mathbf{\Omega}_{\mathcal{M}_{i}\mathcal{F}}(\mathbf{c})e^{\mathrm{inc}}\\ &=\mathbf{\Omega}_{\mathcal{M}\mathcal{M}_{i}}(\mathbf{c})\mathbf{p}_{\mathcal{M}_{i}}(\mathbf{c}).\end{split}$ (31)
@@
Using Weyl’s inequality for singular values [72] [73, Theorem 1] yields $|\sigma_{N_{\mathrm{M}}}-\tilde{\sigma}_{N_{\mathrm{M}}}|\leq\|\mathbf{B}\|_{2}$. Then, using the reverse triangle inequality, we bound $\tilde{\sigma}_{N_{\mathrm{M}}}$ as $\tilde{\sigma}_{N_{\mathrm{M}}}\geq\sigma_{N_{\mathrm{M}}}-\|\mathbf{B}\|_{2}$, provided that $\sigma_{N_{\mathrm{M}}}\geq\|\mathbf{B}\|_{2}$ because otherwise we would have the trivial inequality $\tilde{\sigma}_{N_{\mathrm{M}}}\geq 0$. Finally, defining $\tilde{\mu}\triangleq\|\mathbf{B}\|_{2}.\|\mathbf{A}^{-1}\|_{2}=\frac{\|\mathbf{B}\|_{2}}{\sigma_{N_{\mathrm{M}}}}$ and inverting both sides of the previous inequality yields